The following direct materials and direct labor data pertain to the operations of Laurel Company for the month of August.

Costs

Actual labor rate $15 per hour
Actual materials price $140 per ton
Standard labor rate $14.50 per hour
Standard materials price $142 per ton

Quantities

Actual hours incurred and used 3,800 hours
Actual quantity of materials purchased and used 1,500 tons
Standard hours used 3,850 hours
Standard quantity of materials used 1,490 tons


(a)

Compute the total, price, and quantity variances for materials and labor.

Total materials variance $enter a dollar amount select a type of the variance Neither favorable nor unfavorableFavorableUnfavorable
Materials price variance $enter a dollar amount select a type of the variance UnfavorableNeither favorable nor unfavorableFavorable
Materials quantity variance $enter a dollar amount select a type of the variance FavorableUnfavorableNeither favorable nor unfavorable
Total labor variance $enter a dollar amount select a type of the variance UnfavorableNeither favorable nor unfavorableFavorable
Labor price variance $enter a dollar amount select a type of the variance UnfavorableNeither favorable nor unfavorableFavorable
Labor quantity variance $enter a dollar amount select a type of the variance Neither favorable nor unfavorableFavorableUnfavorable
Answer #1

Total materials variance = 1490*142 - 1500*140 = $1580 Favorable

Materials price variance = (142-140)*1500 = $3000 Favorable

Materials quantity variance = (1490-1500)*142 =$1420 Unfavorable

Total labor variance = 3850*14.50 - 3800*15 = $1175 Unfavorable

Labor price variance = (14.50-15)*3800 = $1900 Unfavorable

Labor quantity variance = (3850-3800)*14.50 = $725 Favorable